# Intertwined Orders and the Physics of High Temperature Superconductors

### v1

@@ VII Microscopic Models of Intertwined Orders
The fact that we have to deal with strongly interacting systems led to the exploration of new approaches such as the large-$N$ and large-$S$ treatments of quantum antiferromagnets by Affleck and Marston [152] and by Read and Sachdev [153, 154, 155]. These novel approaches, which required the generalization of the global symmetry from $SU(2)$ to $SU(N)$ and the consideration of representations higher than the fundamental, led to the discovery of flux phases and valence bond crystals. Wen, Wilczek and Zee showed that such models naturally lead to a possible chiral spin liquid which spontaneously breaks time reversal invariance [156], and that had been conjectured by Kalmeyer and Laughlin [157] using an analogy withe the Laughlin wave function for a fractional quantum Hall state of bosons (see the recent numerical work by Huang, Gong and Sheng [158]). The large-$N$ formulation was also suitable to include the effects of hole-doping and to a mean field theories of d-wave superconductivity, as in the work by Lee, Nagaosa and Wen [159], although requiring the use of somewhat uncontrolled approximations. Recent numerical DMRG studies have confirmed that lightly doped spin liquid ground states are indeed superconductors [160].
@@ Acknowledgements.
My work on this problem has benefitted enormously from the collaboration and insights of many people, both experimentalists and theorists. I am particularly grateful and indebted to Steven Kivelson with who I had the privilege to have shared the development of these ideas which constitute a reformulation of this problem. I have also benefitted from the collaboration with several theorists including Erez Berg, Rodrigo Soto-Garrido and Yuxuan Wang, as well as many experimentalists, particularly John Tranquada, Dale Van Harlingen (who passed away in 2024), Peter Abbamonte, J. C. Séamus Davis, Vidya Madhavan and many others. This work is supported in part by the National Science Foundation grant No. DMR 2225920 at the University of Illinois.
1T-TiSe2 $1T$-TiSe2
BaFeAs122 Ba(Fe1-xCox)2As2
BSCCO Bi2Sr2CaCu2O8+δ
BSNA Ba1-xSrxNi2As2
CDW Charge Density Wave
EuRbFeAs EuRbFe4As4
FF Fulde-Ferrell
FFLO Fulde-Ferrell-Larkin-Ovchinnikov
HTSC High-Temperature Superconductor
LBCO La2-xBaxCuO4
LBCO-1/8 La1.875Ba0.125CuO4
LESCO La1.8-xEu0.2SrxCuO4
LNSCO La1.6-xNd0.4SrxCuO4
LO Larkin-Ovchinnikov
LSCO La2-xSrxCuO4
NbSe2 NbSe2
PDW Pair Density Wave
SDW Spin Density Wave
SC Superconductor
YBCO YBa2Cu3O6+x
UTe2 UTe2

### v2

@@ VII Microscopic Models of Intertwined Orders
[…]
@@ Acknowledgements.
My work on this problem has benefitted enormously from the collaboration and insights of many people, both experimentalists and theorists. I am particularly grateful and indebted to Steven Kivelson with who I had the privilege to have shared the development of these ideas which constitute a reformulation of this problem. I have also benefitted from the collaboration with several theorists including Erez Berg, Rodrigo Soto-Garrido and Yuxuan Wang, as well as many experimentalists, particularly John Tranquada, Dale Van Harlingen (who passed away in 2024), Peter Abbamonte, J. C. Séamus Davis, Vidya Madhavan and many others. My work is supported in part by the National Science Foundation grant No. DMR 2225920 at the University of Illinois.